Back to the NIKA2 main page

Daily Reports

Special NIKA2 Pool on behalf of the DDT E03-20 project, inside a heterodyne Pool (Manager: Mónica Rodríguez). Observing on site. Times referred to LT.

May 28th, Friday

22h15 - 04h30; Ángel: Opacity (tau[225 GHz]) above 0.5, with some spikes up to 0.95 during the shift. First pointing and focus on Mars (low altitude). Beam sizes a bit broadened: ~14.5", 18.5", 14.3"; ToD +/-2 Jy/b, rms ~146 mJy/b (2mm). Focus and pointing on 1226+023 and then on 1156+295 (close to the science target). Monitoring opacity and stability before starting science observations, looking at the beam geometry too. At 23:30 it was decided to start the observations of UDG584 field (E03-20). A total of 7 repetitions (24 min. each) were gathered, starting with fluctuations in time of +/-2 Jy/b at 2 mm and +/-5 Jy/b at 1 mm. Mean rms of ~55 mJy/b at 2 mm. Stopped at 03:20 when the target was close to El.=30 deg. Then a skydip (s29) was tried: too large variation of NIKA2 taus, barely valid. Secondary flux standards MWC349 (s37) and NGC7027 (s39) observed with calib_1scan routine (beam sizes: 12.8", 17.5", 12.9"). EMIR Pool observations continued after 04:30.

May 29th, Saturday

22:45 - 00h00; Angel, Mónica: Mean opacity of about 0.4 throughout the shift, rising to 0.8 in several opportunities. More unstable than yesterday. Pointing and focus on quasars 3C273 and 3C279. Monitoring opacity and stability before calibrations. Best beam sizes of 12.5", 17.8", 12.6" before try a beammap (s315) on 3C273: some instabilities. A new skydip (s316) but under similar conditions to the one obtained on yesterday's shift. Moving to science target at 00:20; relatively stable (Tod +/-2Jy/b at 2 mm). Beam sizes 11.3", 17.6", 11.5". Only one repetition obtained. Instabilities increasing. We decided to continue EMIR Pool observations at ~01:00.

DailyReportsNika2Pool250521 (last edited 2021-06-06 11:50:58 by NikaBolometer)